OpenStudy (anonymous):

I HAVE AN EXAM TOMORROW, PLEASE HELP ME!!! f(x)= 2x+3 what is f^-1 (basically I think they're asking what is the inverse)

OpenStudy (ammarah):

dont u just switch x and y then to find the inverse?

OpenStudy (dtan5457):

Exactly. They are asking for the inverse. Do you know how to do that?

OpenStudy (anonymous):

ok um let me just send you guys the actual question

OpenStudy (dtan5457):

Okie.

OpenStudy (anonymous):

here:

OpenStudy (campbell_st):

you are correct in swapping x and y then you need to make y the subject of the eqution

OpenStudy (anonymous):

I did that, and I got A, not D

OpenStudy (anonymous):

Like if you want a better look at the graphs:

OpenStudy (anonymous):

A:

OpenStudy (anonymous):

B:

OpenStudy (anonymous):

C:

OpenStudy (anonymous):

D:

OpenStudy (anonymous):

E is none of the above

OpenStudy (campbell_st):

well inverse functions are reflected in the line y = x so for f(x) = 2x + 3 the inverse is\[f^{-1}(x) = \frac{x -3}{2}\]

OpenStudy (dtan5457):

If you think it's A. You got the inverse wrong. Let's look at solving the inverse again y=2x+3 switch x and y x=2y+3 2y=x-3 divide by 2 y^-1=(x-3)/2

OpenStudy (anonymous):

ohhhhhhhhhhhhhhh, and again, I AM A FREAKING IDIOT!!!!!!!

OpenStudy (dtan5457):

Now if you graph that...it should be straight foward

OpenStudy (campbell_st):

so the keys, y- intercept -3/2 slope 1/2

OpenStudy (anonymous):

ok i get it now, thanks so much :) :) :)

OpenStudy (dtan5457):

Or if you aren't allowed to use a calculator y=2x+3 Plug random x values and get a bunch of (x,y) coordinates. Inverse is just the coordinates reversed For example If a coordinate for y=2x+3 is -5,-7 the inverse coordinate is -7,-5